OpenStudy (anonymous):

What is the 7th term of the geometric sequence 3, 9, 27, …?

OpenStudy (anonymous):

Here, a = 3 r = 9/3 = 3 So, nth term = ar^(n-1) So, put n = 7, 7th Term = (3)(3)^(7 - 1) So, 7th term will be : 3^7 = 2187 is the answer...

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Welcome dear...